OpenStudy (anonymous):

Hello! Any math tutors in the house? Calculus Q: I'm given a function where x<-2 and then another f(x) where x is greater than or equal to -2. The second f(x) (with the greater or equal to sign) has the variable "a" in it. The question then is: What value must be chosen for a in order to make this function continuous at -2?

OpenStudy (psymon):

A function is continuous at a point if 1. The limit exists 2. The limit as x approaches a is the same answer as f(a)

OpenStudy (anonymous):

so would I substitute (-2) into the functions. set both functions equal to eachother and solve for a?

OpenStudy (psymon):

Do you have the actual functions?

OpenStudy (anonymous):

yup. this is my first time on this site. I just don't know how to type them in here, but i'll try

OpenStudy (psymon):

Ah. Well, you can hit the equation button below the typing box. Feel free to get used to it.

OpenStudy (anonymous):

oh! I never noticed that! are you a university student as well?

OpenStudy (psymon):

Yeah, I am. Im actually a math tutor where I go to school, lol. Im just the.....least experienced math tutor, but meh xD

OpenStudy (anonymous):

ahhhhh. I hope you're getting paid for this.

OpenStudy (psymon):

Not for on this site. I just come onhere not only to hang out, but doing this helps me practice and make sure I know what I need to know or find out what I need to learn.

OpenStudy (psymon):

To be honest with you, a month ago I probably couldnt have given you that definition I gave. I intuitively knew it, but wouldnt have known the actual definition of continuity.

OpenStudy (e.mccormick):

\(\begin{cases} f_1(x) & \text{ if } x<-2 \\ f_2(x) & \text{ if } x\ge -2 \end{cases}\) Right click to see the \(\LaTeX\) code used to do cases.

OpenStudy (anonymous):

That's pretty impressive. To be honest, I hate math. I need this course to apply for graduate school. Yeah, i'm taking a first year calculus course at ubc, it's been quite the roller coaster for me.

OpenStudy (anonymous):

k, ill post the equation on here, maybe someone can help. when i equate them to eachother, one of the denominator equals 0, which i know is no no.

OpenStudy (psymon):

Well, just to see what your two pieces to the function are mighthelp.

OpenStudy (anonymous):

f(x) = 6x\[f(x)=6x^3 + 13x^2 + 12x + 20 / x + 2\]

OpenStudy (anonymous):

the first one doesn't count lol

OpenStudy (anonymous):

i'll write the second one now

OpenStudy (psymon):

Lol, alright.

OpenStudy (anonymous):

\[f(x) -6x^2 + 6x + a\]

OpenStudy (e.mccormick):

\(f(x)=\dfrac{6x^3 + 13x^2 + 12x + 20}{x + 2}\) perhaps?

OpenStudy (anonymous):

for the first one: x is less than -2. for the second one x is greater than or equal to -2.

OpenStudy (psymon):

Yeah, so seems like youre looking for a value of a such thatthe two graphs intersect at x = -2

OpenStudy (anonymous):

oooo thanks. that's fancy.

OpenStudy (psymon):

Yeah, theres a fraction button, too xD

OpenStudy (anonymous):

q: What value must be chosen for a in order to make this function continuous at -2? ... why can't i just equal them to each other and sole for a

OpenStudy (e.mccormick):

\(f(x)=\begin{cases} \dfrac{6x^3 + 13x^2 + 12x + 20}{x + 2} & \text{ if } x<-2 \\ \\ -6x^2 + 6x + a & \text{ if } x\ge -2 \end{cases}\)

OpenStudy (anonymous):

omgd. e.mccormick. you're amazing.

OpenStudy (e.mccormick):

Nah, I have just use \(\LaTeX\) longer than many here and can hand type it. I have done some papers with it.

OpenStudy (psymon):

You can actually factor that top piecewise to get the x+2 to cancel out. Will make finding out your y-value much easier.

OpenStudy (anonymous):

but x is cubed. and 20 doesn't have an x.

OpenStudy (e.mccormick):

Yah, you basically need to find what it would be at -2 on the top if it was valid. Then you find a value of the bottom that would have that same answer.

OpenStudy (anonymous):

yeah, just need to get rid of the denominator

OpenStudy (psymon):

I tested it just to see. Divide it out and check for yourself, it does divide well :3

OpenStudy (anonymous):

kk, i'll give it a shot, thanks! xoxo

OpenStudy (psymon):

Yep, sure. Once you divide it out, you can plug in x = -2 for that function and find the y-value. Then your job is to fine a such that x = -2 in that function produces the same y-value.

OpenStudy (anonymous):

thanks!!! i'll post an update if i get it. lol

OpenStudy (anonymous):

i mean, "when" lol

OpenStudy (psymon):

Yeah, sure xD

OpenStudy (e.mccormick):

I'd show it to you on a graph, but that would give it away.

OpenStudy (anonymous):

wow. I can't seem to factor the x cubed. any suggestions?

OpenStudy (anonymous):

i know i have to factor out (x+2) from each term

OpenStudy (anonymous):

atleast i think

OpenStudy (psymon):

Well, I just literally divided it by x+2 to check 6x^2 + x + 10 x+2 | 6x^3 + 13x^2 + 12x + 20 -[6x^3 +12x^2] x^2 + 12x -[ x^2 + 2x] 10x + 20 -[10x + 20] 0

OpenStudy (anonymous):

i want to curl up in a corner and never wake up.

OpenStudy (psymon):

So if you did the division, obviously (x+2) is a factor because of no remainder. What I have left is what happens after that (x+2) term cancels. So since the x+2 term in the denominator is now gone, we can plug in x = -2 to see what the y-value is: 6(-2)^2 -2 + 10 = 32 So now we need a value of a in the other function such that as x approaches -2, the answer is 32. And no o.o No need to beat yourself up or something o.o

OpenStudy (e.mccormick):

Did you go over synthetic division? There is also factor by grouping as a possibility.

OpenStudy (anonymous):

no, I just googled it. grouping I learned a couple of years ago.

OpenStudy (e.mccormick):

http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

a=68!

OpenStudy (anonymous):

holy shinizzle.

OpenStudy (psymon):

Lol.

OpenStudy (anonymous):

thanks for your help Psymon and e.mcc (going to call you mc for short)

OpenStudy (e.mccormick):

FYI: Calculus requires and uses every type of math before it. For example, take a very simple integral: \(\int 2 x \text{d}x \) Once you learn the rules, you know this has a value added to the exponent and the result of that is divided out of the whole thing. So: \(\int 2 x \text{d}x \implies \int 2 x^1 \text{d}x \implies \dfrac{2x^{1+1}}{2}+c\implies x^2+c\) Now, I just used \(1+1=2\) and \(\dfrac{2}{2}=1\) and was doing calculus.

OpenStudy (e.mccormick):

And here is that graph I talked about: https://www.desmos.com/calculator/na0jflqgdz Very good tool to use for checking yourself!

OpenStudy (anonymous):

Thanks you two. appreciate it MC!

OpenStudy (psymon):

Yep yep.

OpenStudy (e.mccormick):

And a great site for brushing up on all the math you forgot: https://www.khanacademy.org/

OpenStudy (e.mccormick):

I used KA to get back up to speed and it made a HUGE difference! Have fun!

OpenStudy (anonymous):

oh man. that site is a life saver. esp for logs. need to brush up on some algebra. this girl is pretty amazing too: http://www.youtube.com/user/TheIntegralCALC

OpenStudy (e.mccormick):

Yah! I saw a couple of hers. Good. Ever seen PatricJMT's videos?

OpenStudy (e.mccormick):

PatrickJMT that is...

OpenStudy (anonymous):

nope on youtube?

OpenStudy (e.mccormick):

Yah. He lists his by topic and is a math major. Does it on paper or white board woth a marker, so all you see is the math.

OpenStudy (anonymous):

perfect. wrote it down on my notebook.

OpenStudy (anonymous):

are you from north america as well?

OpenStudy (e.mccormick):

California. You are a tad furthur north in Canada.

OpenStudy (anonymous):

I've always wanted to explore California. Your summers seem amazing!

OpenStudy (e.mccormick):

Well, it depends on where. The state ranges from snow country in the winter to, well, dry and sunny in the winter. In the summer that becomes sunny and breezy to roasted.

OpenStudy (anonymous):

eeeeek. well, i've always wanted to visit there. esp after the hit song california girls. jk jk

OpenStudy (e.mccormick):

Just don't visit this part in the middle of summer: http://www.livescience.com/38054-why-death-valley-hot.html

OpenStudy (anonymous):

NEVER. that's our future billions of years from today.